# Leveraging Diffusion Models for Parameterized Quantum Circuit Generation

### v2

@@ IV-B GHZ Task:
Figure 2: Comparison of fidelity distributions and evaluation metrics for two different gate sets: gs1 (top) and gs2 (middle) and generalization ability of our DM on 4 qubits for gs1 (bottom). Shown are the fidelity scores of synthesized PQCs under varying guidance scales (left) and their respective evaluation results, e.g. invalid circuits and uniqueness (right).
In this section, we present the results obtained by applying our DM to the GHZ state-preparation task under two distinct gate sets, namely $gs1$ and $gs2$. For each task, we synthesize 500 samples and 250 for 4-qubits on $gs1$, thereby ensuring statistically robust insights. As illustrated in Fig. 2, we analyze key metrics, including fidelity distributions, error counts, structural uniqueness, and parameter uniqueness. Here, parameter uniqueness refers to the diversity in continuous gate parameters, distinguishing otherwise topologically identical circuits, thereby enabling distinct quantum operations. Additionally, we discuss the critical role of  CFG in modulating circuit synthesis, leading to increasingly accurate and diverse PQCs.
For $gs1$ Fig. 2 (top) our DM consistently achieves near-perfect fidelity on three-qubit GHZ circuits, effectively converting upwards of 95% of its initial continuous tensors into valid, single-gate-per-timestep circuits. The introduction of moderate guidance scales (e.g., 10) pushes fidelity distributions toward the ideal value of $1.0$. Crucially, even at perfect fidelity targets, the model produces diverse circuit topologies. This structural variability is invaluable for practical considerations such as noise-resilience testing and hardware calibration.
An equally important observation is the model’s capacity to produce structurally diverse high-fidelity circuits. Even when tasked with the same target fidelity, the DM naturally explores multiple circuit topologies, demonstrating a talent for balancing performance with variability that could be beneficial in tasks like noise-robustness analysis or hardware calibration. Although trained solely on three-qubit examples, the DM exhibits promising scalability to four- and five-qubit scenarios Fig. 2 (bottom). At four qubits, around 30% of generated circuits reach perfect fidelity, but with a significant higher error count. At five qubits, only a few circuits achieve the ideal GHZ state, and the error count grows substantially. Furthermore, our results indicate promising scalability. The overall computational overhead remains low, enabling the rapid generation of large circuit ensembles and facilitating efficient selection of high-fidelity designs. This capability to generalize beyond the qubit counts seen in training underscores the fundamental adaptability of diffusion-based generative models. Nevertheless, while out-of-distribution fidelity declines as qubit count increases, these findings motivate further refinements—including curriculum learning over multi-qubit data or advanced sampling strategies to bridge performance gaps in higher-dimensional spaces. By systematically enhancing the training pipeline, we anticipate even more robust GHZ synthesis at larger qubit scales.
Compared to $gs1$, the $gs2$ alphabet incorporates hardware-native gates, such as ’sx’ and ’rzz’, complicating the synthesis task. As shown in Fig. 2 (middle), the absence of CFG initially yields circuits spread across lower fidelity values, reflecting the increased difficulty of orchestrating these more complex building blocks. However, once CFG is introduced, fidelity improves dramatically. At a guidance scale of 5, the DM already reaches high fidelity in mostof circuits though with 50 higher error count and less unqiues in structure. Scaling guidance to 15 pushes fidelity even further, increases the accuracy. This jump demonstrates how CFG can effectively counterbalance the tougher hardware constraints and high-dimensional parameter spaces inherent in $gs2$.
A particularly notable discovery is the emergence of layer-like circuit motifs, where single-qubit rotations precede entangling gates, illustrated in Fig. 3. The DM autonomously learns these structural templates, apparently having distilled key architectural heuristics from the training data. Importantly, each “layer” is composed of parameterized gates with unique angle settings, illustrating that diversity is not merely confined to gate ordering but also extends into the continuous parameter space.
The strong performance of our DM in synthesizing hardware-native circuits suggests a promising pathway for practical, device-level quantum compilation. By adjusting guidance scales and training on broader gate sets, one could systematically adapt the pipeline to specific backend constraints, which is an important step forward in bridging the gap between algorithmic design and near-term quantum hardware.
Figure 3: Representation of two example circuits generated by our diffusion model illustrating how the model arranges multi-layer gate structures for GHZ tasks. synthesized circuit from $gs2$ with high fidelity guidance (left) and a circuit intended for QML classification with high accuracy guidance (right).
Together, these results showcase the DM’s notable adaptability and scalability, demonstrating strong performance under stricter device constraints. Even for gate sets that demand more intricate strategies, the diffusion model augmented by CFG continues to discover high-fidelity, structurally sophisticated circuits, firmly establishing its promise for automated, hardware-aware quantum circuit synthesis.
@@ IV-C ML task
Figure 4: Output distribution to match different accuracy targets demonstrating the models adaptability and controlled circuit generatio: Synthesized PQCs with target accuracy of 0.2 (left) and synthesized PQCs with target 0.7.
Beyond GHZ-state preparation, our diffusion model demonstrates promising potential in  QML classification tasks. To evaluate its performance, we generated 250 circuits explicitly conditioned on distinct accuracy targets (e.g., 0.7 or 0.2), assessing the model’s ability to synthesize diverse PQCs that align closely with user-defined performance levels. Notably, the dataset posed additional challenges in achieving extremely high accuracy, as it contained fewer circuits with accuracy scores of 0.9 or above. Nevertheless, the model successfully guided circuit generation toward intermediate accuracy values, effectively producing circuits centered around target accuracies such as 0.7 Fig.4 (left) and 0.2 Fig.4 (right). Leveraging CFG, our approach demonstrates the capacity to systematically explore a broad range of PQC designs, enabling precise control and rapid generation of task-oriented quantum classifiers without the need for extensive retraining or iterative simulations.
Crucially, even PQCs targeting lower accuracies maintain structurally unique designs, emphasizing the model’s extensive exploration capability. This controlled generation capability significantly advances automated PQC design, providing users a powerful tool to systematically explore circuit architectures tailored to specific quantum machine learning tasks. The diffusion model thus establishes a computationally efficient and versatile approach for discovering novel and practical quantum classifier solutions, greatly expanding the landscape of achievable quantum computational functionalities.
Taken together, these outcomes highlight a pivotal advancement in automated PQC design for both state-preparation and QML objectives. Whether aiming for high-fidelity GHZ circuits or precisely targeted classifier accuracies, the diffusion model - amplified by CFG - delivers robust, controllable, and scalable circuit synthesis. Its generalization capability and low computational overhead further solidify its practicality, as users can swiftly generate large ensembles of candidate solutions and select those that best suit their particular hardware or application constraints. Such versatility not only enriches quantum algorithm design workflows but also marks an important step toward broader adoption of generative techniques in quantum computing research and industry.
@@ V Discussion
Our findings confirm that diffusion models can effectively synthesize high-quality PQCs without costly quantum-mechanical simulations or iterative parameter optimization. Not only does the model consistently meet or surpass the benchmarks of traditional generative methods, but it does so while maintaining considerable adaptability. As demonstrated in our experiments, CFG emerges as a crucial factor, enabling users to smoothly adjust fidelity or accuracy targets and quickly explore the resulting diverse architectural landscapes. Moreover, by simply editing textual prompts or gating configurations, the model accommodates a broad range of hardware constraints and design preferences, mirroring the flexibility of diffusion-based systems in other domains. This inherent malleability paves the way for advanced editing operations such as circuit masking or partial design overwrites, extending well beyond the usual boundaries of classical search or reinforcement learning approaches. Altogether, these attributes suggest that diffusion models not only excel in near-term tasks like GHZ-state generation or quantum classification but may also serve as robust, general-purpose frameworks for a variety of quantum synthesis challenges.
@@ VI Conclusion and Outlook
In summary, we have introduced and evaluated a diffusion-based pipeline for generating parameterized quantum circuits across a diverse array of tasks and hardware settings. Our empirical results underscore its capacity to deliver high-fidelity GHZ states, optimize QML classifiers, and flexibly integrate hardware-related design constraints. Crucially, the method requires no iterative parameter fitting or direct quantum-state simulation during generation, offering a scalable, data-driven alternative to conventional synthesis strategies.
Looking ahead, future work could involve deeper exploration of multi-qubit scaling, integration of noise-mitigation strategies, and expanded application to broader classes of variational quantum algorithms. The apparent synergy between classifier-free guidance and diffusion modeling also opens new avenues for task-specific customizations—enabling nuanced trade-offs between circuit depth, fidelity, and hardware connectivity. We anticipate that these developments will further consolidate diffusion-based models as a cornerstone in automated quantum circuit design, driving innovation in both academic and industrial quantum research.

### v3

@@ IV-B GHZ Task
Figure 2: Comparison of fidelity distributions and evaluation metrics for two different gate sets: gs1 (top) and gs2 (middle) and generalization ability of our DM on 4 qubits for gs1 (bottom). Shown are the fidelity scores of synthesized PQCs under varying guidance scales (left) and their respective evaluation results, e.g. error counts (invalid circuits) and uniqueness (right).
In this section, the results of applying our diffusion model to the GHZ state-preparation task with two distinct gate sets, $gs1$ and $gs2$, are presented. For each setting, 500 samples were synthesized to ensure statistically meaningful conclusions. As shown in Fig. 2, several key metrics are analyzed: the fidelity distributions (left), as well as evaluation metrics including error counts, structural uniqueness, and parameter uniqueness (unique hash) (right). Here, error counts denote the number of outputs from the diffusion model that could not be decoded into valid quantum circuits. Structural uniqueness quantifies the number of structurally distinct circuit architectures generated, while parameter uniqueness measures the diversity of continuous gate parameters among circuits with identical architecture, capturing variations that yield distinct quantum operations. Uniqueness metrics were evaluated only for circuits that reached the target performance.
Figure 3: Output distribution to match different accuracy targets demonstrating the model’s adaptability and controlled circuit generation: synthesized PQCs with target accuracy of $0.2$ (left) and synthesized PQCs with target $0.8$ (right).
In the following, the critical impact of CFG on circuit synthesis is discussed, demonstrating how adjusting guidance parameters modulates the generation of PQCs tailored to the desired task. For gate set $gs1$ (Fig. 2, top), and with a guidance scale of 00, the diffusion model generates only valid PQCs, each featuring a single gate per timestep and no decoding errors. The resulting fidelity distribution peaks around $0.5$. Increasing the guidance to $10$ shifts the fidelity sharply toward the target value of $1.0$, with nearly all synthesized circuits achieving perfect fidelity. Notably, even at this high fidelity, the model continues to produce a diverse array of circuit architectures; the majority of circuits exhibit structural uniqueness, and all demonstrate parameter (hash) uniqueness. This structural diversity is particularly valuable for applications requiring noise robustness or device calibration, as it enables exploration of multiple solutions. In contrast, the $gs2$ gate set (Fig. 2, middle) introduces hardware-native gates like ‘sx’ and ‘rzz’, making the synthesis task more challenging. Without CFG, the model produces circuits with fidelities spread across lower values, reflecting the increased difficulty of synthesizing with a more complex alphabet. Introducing guidance modestly increases fidelity, but the benefit plateaus quickly, reaching a maximum at a guidance scale of $2.5$. Beyond this, further increases in guidance lead to a substantial drop in both structural and parameter uniqueness, while error counts rise sharply. This highlights the added complexity and trade-offs associated with hardware-constrained synthesis tasks.
When evaluating the model’s generalization ability in a zero-shot setting (Fig. 2, bottom), it is observed that a model trained exclusively on 3-qubit circuits is nevertheless able to generate valid 4-qubit circuits, despite never having encountered such configurations during training. Moreover, increasing the guidance scale enables the model to shift the fidelity distribution toward the desired target value, displaying that the DM understands how PQCs must be structured and parameterized to achieve a given task. However, this comes at the cost of a significantly increased error rate, suggesting that finetuning the model on multi-qubit data could further enhance performance. For five-qubit circuits, the diffusion model is able to produce a small number of circuits that achieve the ideal GHZ state when guidance is applied, but the error count rises substantially with increasing circuit size. These findings underscore both the scalability potential and current limitations of the approach, highlighting the value of targeted finetuning for more challenging generalization tasks.
@@ IV-C ML task
Beyond GHZ-state preparation, the diffusion model shows strong performance on QML classification tasks. To assess this, 250 circuits were generated with explicit conditioning on various accuracy targets, allowing evaluation of the model’s ability to produce PQCs that meet specific performance requirements. Although the dataset included few circuits achieving high accuracy ($\geq$ 0.7), the model reliably generated circuits clustered around the requested accuracy targets, as seen for $0.2$ (Fig. 3, left) and $0.8$ (Fig. 3, right), as guidance increased, while maintaining high diversity in both circuit structures and parameter values, along with a low error rate. This highlights the adaptability of the model, as it successfully transitions from generating circuits for fixed-state objectives to producing circuits that perform well on classification tasks using linearly separable data.
IV-D Inference Efficiency and Structural Insights
Figure 4: Example PQCs with layerwise gate structures generated by DM: Synthesized PQC from $gs2$ with high fidelity guidance (left) and a classification PQC with high accuracy guidance (right).
A significant discovery within this context is the autonomous emergence of structured, layer-like circuit ansätze composed of alternating single-qubit rotations and entangling gates (Fig. 4). The consistent appearance of these patterns suggests that the model has implicitly distilled and generalized fundamental architectural principles from the training data. Importantly, this structural regularity does not limit parameter diversity, as each layer maintains distinct continuous parameter values. To further assess the practicality and scalability of our model, which was trained only on 3-qubit circuits with a maximum of 16 gates for the $gs1$ task, we systematically analyzed the model’s inference efficiency and output quality as a function of circuit size and gate count. Tables I and II summarize key results. Table I shows the impact of increasing the number of qubits (with fixed gate count), where generation and conversion times (time needed to decode the tensor into a circuit) remain almost constant, but higher qubit counts lead to reduced numbers of synthesized high-fidelity circuits, less structural and parameter uniqueness, and a sharp rise in decoding errors. In contrast, Table II varies gate count (for three qubits) and reveals that the model produces the highest number of unique, target-fidelity circuits at intermediate depths (12–24 gates), while both shallow and very deep circuits show increased error rates and lower uniqueness. Together, these tables illustrate the model’s strong inference efficiency, but also highlight the practical trade-offs between scalability, diversity, and reliability in quantum circuit synthesis.
TABLE I: Scalability and Inference Latency over Qubit Number for $gs1$ (Max Gates = 16, Guidance = 10)
Qubit Count Gen. Time (s) Conv. Time (s) High Fid. Count Uniq. Struc. Uniq. Hash Error Count
3 54.578 1.471 82 82 82 0
4 56.221 0.739 23 23 23 51
5 56.555 0.192 2 2 2 88
TABLE II: Scalability and Inference Latency over gate count for $gs1$ (Qubit Number = 3, Guidance = 10)
Max. Gates Gen. Time (s) Conv. Time (s) High Fid. Count Uniq. Struc. Uniq. Hash Error Count
8 41.486 1.733 60 60 60 4
12 43.195 1.362 85 80 85 1
16 54.578 1.471 82 82 82 0
20 61.176 1.539 80 79 80 0
24 61.756 1.419 79 79 79 1
28 88.816 1.246 67 67 67 15
@@ V Discussion
In this work, we have demonstrated the potential of DMs as a flexible, scalable, and data-driven approach for synthesizing PQCs. By effectively addressing tasks in quantum state preparation, specifically GHZ state synthesis, as well as QML classification, our approach potentially offers significant advantages over traditional circuit synthesis methods. The experimental results confirm that diffusion-based models not only consistently synthesize valid circuits meeting targeted fidelity and accuracy (Fig. 2 and Fig. 3), but it also exhibits strong zero-shot generalization, as the model is able to generate PQCs that achieve the desired targets even for circuit sizes not seen during training (Fig. 2, bottom), which indicates a genuine understanding of how circuits must be structured and parameterized to fulfill specific objectives. Moreover, the approach offers practical control over the synthesis process through CFG, which enables users to navigate the trade-off between performance and diversity: higher guidance values promote precise, target-oriented circuit generation but may reduce diversity and error count, whereas lower guidance facilitates greater diversity in circuit generation, a valuable property in scenarios that require noise-tolerant designs or thorough calibration. Beyond these performance aspects, our approach autonomously uncovers fundamental architectural heuristics, such as layer-like ansatz structures, consisting of single-qubit rotations followed by entangling gates (Fig. 4). The ability of the model to distill and reproduce these classical patterns illustrates a holistic exploration of the PQC solution space, going beyond mere interpolation between training samples.
A key practical strength of the diffusion-based approach lies in its computational efficiency during inference. The circuit generation time remains nearly constant as qubit number increases and scales slightly with gate count (Tables III). This strong scalability avoids the common bottleneck with search-based approaches, which typically require costly quantum simulations within optimization loops [12, 4, 5, 26, 2]. However, this efficiency entails a trade-off: as system size and circuit depth grow, the quantity of high-fidelity, structurally unique circuits decline, accompanied by increased error rates. Such behavior underscores the quantum nature of the underlying problem, wherein scaling to larger circuits remains inherently challenging due to exponential growth in complexity, which can be avoided with finetuning on bigger qubit number or curriculum learning.
While the current results are promising, several limitations persist. When generalizing to larger systems or more complex gate sets, the model encounters increased invalid outputs and decoding errors, reflecting the intrinsic complexity introduced by hardware-native constraints. The generation of circuits under restrictive hardware conditions is particularly challenging and may demand further model enhancements, broader dataset curation, or explicit integration of hardware-noise models within the training process. Moreover, assembling suitable training datasets for large-scale models, especially circuits requiring classical unitary computations, poses considerable challenges, highlighting the need for alternative conditioning methods, such as employing Hamiltonian-based representations for unitary evolutions, to improve efficiency.
@@ VI Conclusion and Outlook
In summary, we have introduced and evaluated a diffusion-based pipeline for generating PQCs across diverse tasks and hardware gate sets. Our empirical results underscore its capacity to deliver high-fidelity GHZ states, optimize classifiers, and flexibly integrate hardware-related design constraints. Crucially, the method requires no iterative parameter optimization or direct quantum-state simulation during training and generation, offering a scalable alternative to existing synthesis strategies.
Looking ahead, several directions remain for further development. Expanding the model to handle multi-parameter gates and leveraging starting ansätze tailored to specific quantum tasks could enhance both expressiveness and synthesis efficiency. Further research could explore larger qubit systems, integrate noise mitigation, and adapt the approach to a wider range of variational quantum algorithms. The demonstrated synergy between CFG and diffusion modeling presents opportunities for task-specific customization, including nuanced trade-offs between circuit depth, fidelity, and hardware constraints. Collectively, these advancements could further consolidate DMs as a foundation for automated quantum circuit design and accelerate progress across both academic and industrial quantum research.